# Superconductivity from spin-canting fluctuations in rhombohedral graphene

### v3

@@ Appendix A Calculation of spin susceptibilities
In the last line of Eqs. (47) and (48) we have performed an expansion to leading order in frequency and momentum. Note that while $\chi_{1}(i\nu,{\boldsymbol{q}})$ is an even function of its arguments, $\chi_{2}(i\nu,{\boldsymbol{q}})$ is an odd function. We further expect the absence of any linear-in-${\boldsymbol{q}}$ term in $\chi_{2}$ due to the model’s rotation symmetry, together with analyticity of $\chi_{2}(q)$; the latter property precludes a $\sim|{\boldsymbol{q}}|$ contribution.
The expression for the parameter $D$ is complicated and depends on the detail of band dispersion, so we do not write it down explicitly. On the other hand, the parameters $C_{0}$, $\kappa$ and $\gamma$ can be computed straightforwardly from the ${\boldsymbol{q}}=0$ limit of the above expressions:
@@
$C_{0}=\chi_{1}(0,0)=\frac{n_{0}}{\bar{m}}=\frac{1}{V+J},\quad\kappa=\frac{n_{0}}{4\bar{m}^{3}}=\frac{1}{4(V+J)\bar{m}^{2}},\quad\gamma=\frac{n_{0}}{2\bar{m}^{2}}=\frac{1}{2(V+J)\bar{m}}=2\kappa\bar{m}.$ (51)
Above we used the saddle-point condition from Eq. (15) to simplify $C_{0}$, $\kappa$ and $\gamma$, leading to the expressions quoted in the main text.
@@ C.1 Magnon spectrum: order of magnitude estimates and numerics
In this Section we provide estimates of the various scales involved in the magnon spectrum obtained from the poles of the magnon propagators $\mathcal{G}_{b\eta}(q)$ in Eq. (27) of the main text. For simplicity we first consider the critical point $\theta_{0}=0$, where the four magnon modes coalesce into only two solutions due to the restored U(1) spin rotation symmetry: a gapped ($+$) and a gapless, linearly dispersing (--) branch:
@@
For rhombohedral multilayers (in particular following spin-orbit proximitized trilayer experiments [44]), we can estimate $J\sim 200$ meV nm2 and polarization densities (per valley) $n_{0}\sim 2\times 10^{11}$ cm2 relevant for the region where SC is observed. The strength of intra-valley interactions $V$ (which are momentum-independent within our microscopic treatment, Eq. (5)) can be roughly estimated from the value of the gate-screened Coulomb potential
@@
at the Fermi momentum in the relevant low-density regime, $|{\boldsymbol{q}}|\sim k_{F}\sim\left(0.05-0.1\right)a_{0}^{-1}$. Here $q_{e}$ is the electron charge, $\epsilon_{0}$ the vaccum permittivity and $d$ denotes the gate distance, which is of order of a few tens of nanometers in typical experiments, such that $k_{F}d\gg 1$. The estimate for $V$ also depends on the amount of electronic screening in the system, which is difficult to evaluate accurately. In self-consistent Hartree-Fock treatments (e.g. [68, 52, 53, 54]) where screening is accounted for phenomenologically by treating the relative permittivity $\epsilon_{r}$ as a free parameter, a best fit to the experimentally observed phase boundaries [3, 5, 40, 44] is obtained in the range $\epsilon_{r}\sim 20-30$. With these various uncertainties in mind, we estimate $V(|{\boldsymbol{q}}|\sim k_{F})\sim 1-3$ eV nm2. The corresponding ratio of intra-valley to inter-valley interactions in our theory should then be $V/J\sim 5-15$.
@@
We plot the magnon spectrum in Fig. 3 for $D=1$ and $10$ eV-1, and two values of $V/J=5,10$, both in the ranges justified above. The separation of energy scales between the high-energy and the soft magnon modes is quite large for momentum transfers of the order $k_{0}a_{0}\sim 0.01-0.05$, which are expected to be most important for interband pairing that scatter electrons between majority and minority Fermi surfaces. (Note that $k_{0}$ denotes the momentum transfer between the majority and minority Fermi pockets, which can be much smaller than $k_{F}$ especially in the presence of trigonal warping).
@@ Appendix D Derivation of magnon-mediated interaction
$\mathcal{M}_{\xi_{1}\xi_{2}}=\frac{1}{2}\frac{1}{Q^{\theta}_{s}Q^{\phi}_{s}+\gamma^{2}\nu_{n}^{2}}\left(\begin{matrix}Q^{\phi}_{s}&-\gamma\nu_{n}\\ +\gamma\nu_{n}&Q^{\theta}_{s}\end{matrix}\right)+\frac{1}{2}\tau_{\xi_{1}}\tau_{\xi_{2}}\frac{1}{Q^{\theta}_{a}Q^{\phi}_{a}+\gamma^{2}\nu_{n}^{2}}\left(\begin{matrix}Q^{\phi}_{a}&-\gamma\nu_{n}\\ +\gamma\nu_{n}&Q^{\theta}_{a}\end{matrix}\right)$ (72)
@@
$g^{+-,+-}_{\xi_{1},\xi_{2}}(q)$ $=-\frac{1}{2}\frac{Q^{\phi}_{s}+(-i)(-\gamma\nu_{n})+(-i)(\gamma\nu_{n})+(-i)^{2}Q^{\theta}_{s}}{Q^{\theta}_{s}Q^{\phi}_{s}+\gamma^{2}\nu_{n}^{2}}-\tau_{\xi_{1}}\tau_{\xi_{2}}\frac{1}{2}\frac{Q^{\phi}_{a}+(-i)(-\gamma\nu_{n})+(-i)(\gamma\nu_{n})+(-i)^{2}Q^{\theta}_{a}}{Q^{\theta}_{a}Q^{\phi}_{a}+\gamma^{2}\nu_{n}^{2}},$
$=-\frac{1}{2}\frac{Q^{\phi}_{s}-Q^{\theta}_{s}}{Q^{\theta}_{s}Q^{\phi}_{s}+\gamma^{2}\nu_{n}^{2}}-\tau_{\xi_{1}}\tau_{\xi_{2}}\frac{1}{2}\frac{Q^{\phi}_{a}-Q^{\theta}_{a}}{Q^{\theta}_{a}Q^{\phi}_{a}+\gamma^{2}\nu_{n}^{2}},$ (73)
$g^{-+,+-}_{\xi_{1},\xi_{2}}(q)$ $=-\frac{1}{2}\frac{Q^{\phi}_{s}+(-i)(-\gamma\nu_{n})+i(\gamma\nu_{n})+i(-i)Q^{\theta}_{s}}{Q^{\theta}_{s}Q^{\phi}_{s}+\gamma^{2}\nu_{n}^{2}}-\tau_{\xi_{1}}\tau_{\xi_{2}}\frac{1}{2}\frac{Q^{\phi}_{a}+(-i)(-\gamma\nu_{n})+i(\gamma\nu_{n})+i(-i)Q^{\theta}_{a}}{Q^{\theta}_{a}Q^{\phi}_{a}+\gamma^{2}\nu_{n}^{2}},$
$=-\frac{1}{2}\frac{Q^{\phi}_{s}+Q^{\theta}_{s}+2i\gamma\nu_{n}}{Q^{\theta}_{s}Q^{\phi}_{s}+\gamma^{2}\nu_{n}^{2}}-\tau_{\xi_{1}}\tau_{\xi_{2}}\frac{1}{2}\frac{Q^{\phi}_{a}+Q^{\theta}_{a}+2i\gamma\nu_{n}}{Q^{\theta}_{a}Q^{\phi}_{a}+\gamma^{2}\nu_{n}^{2}}$ (74)
@@
$Q^{\theta}_{s}=2\tilde{J}\cos^{2}\theta_{0}+\kappa\nu_{n}^{2}+D{{\boldsymbol{q}}}^{2},\quad Q^{\phi}_{s}=\kappa\nu_{n}^{2}+D{{\boldsymbol{q}}}^{2},$ (75)
$Q^{\theta}_{a}=2\tilde{J}\sin^{2}\theta_{0}+\kappa\nu_{n}^{2}+D{{\boldsymbol{q}}}^{2},\quad Q^{\phi}_{a}=2\tilde{J}+\kappa\nu_{n}^{2}+D{{\boldsymbol{q}}}^{2}.$ (76)
@@
$g^{+-,+-}_{\xi_{1},\xi_{2}}(q)\sim$ $\frac{\tilde{J}\cos^{2}\theta_{0}}{\left[2\tilde{J}\kappa\cos^{2}\theta_{0}+\gamma^{2}\right]\nu_{n}^{2}+2\tilde{J}\kappa\cos^{2}\theta_{0}D{{\boldsymbol{q}}}^{2}}$
$-\tau_{\xi_{1}}\tau_{\xi_{2}}\frac{\tilde{J}\cos^{2}\theta_{0}}{4\tilde{J}^{2}\sin^{2}\theta_{0}+\left[2\tilde{J}\kappa\left(1+\sin^{2}\theta_{0}\right)+\gamma^{2}\right]\nu_{n}^{2}+2\tilde{J}\kappa(1+\sin^{2}\theta_{0})D{{\boldsymbol{q}}}^{2}},$ (77)
$g^{-+,+-}_{\xi_{1},\xi_{2}}(q)\sim$ $-\frac{\tilde{J}\cos^{2}\theta_{0}+\kappa\nu^{2}+D{{\boldsymbol{q}}}^{2}+i\gamma\nu}{\left[2\tilde{J}\kappa\cos^{2}\theta_{0}+\gamma^{2}\right]\nu_{n}^{2}+2\tilde{J}\kappa\cos^{2}\theta_{0}D{{\boldsymbol{q}}}^{2}}$
$-\tau_{\xi_{1}}\tau_{\xi_{2}}\frac{\tilde{J}(1+\sin^{2}\theta_{0})+\kappa\nu^{2}+D{{\boldsymbol{q}}}^{2}+i\gamma\nu}{4\tilde{J}^{2}\sin^{2}\theta_{0}+\left[2\tilde{J}\kappa\left(1+\sin^{2}\theta_{0}\right)+\gamma^{2}\right]\nu_{n}^{2}+2\tilde{J}\kappa(1+\sin^{2}\theta_{0})D{{\boldsymbol{q}}}^{2}},$ (78)

### v4

@@ Appendix A Calculation of spin susceptibilities
In the last line of Eqs. (48) and (49) we have performed an expansion to leading order in frequency and momentum (i.e., when the corresponding terms are small compared to $\bar{m})$. While $\chi_{1}(i\nu,{\boldsymbol{q}})$ is an even function of its arguments, $\chi_{2}(i\nu,{\boldsymbol{q}})$ is an odd function. We further expect the absence of any linear-in-${\boldsymbol{q}}$ term in $\chi_{2}$ due to the model’s rotation symmetry, together with analyticity of $\chi_{2}(q)$; the latter property precludes a $\sim|{\boldsymbol{q}}|$ contribution. We note that $\chi_{1}$ will inherit a non-analytical term $\sim|\nu|$ for $|{\boldsymbol{q}}|>k_{F}^{+}-k_{F}^{-}$ due to Landau damping, as detailed in Appendix F.
The expression for the parameter $D$ is complicated and depends on the detail of band dispersion, so we do not write it down explicitly. On the other hand, the parameters $C$, $\kappa$ and $\gamma$ can be computed straightforwardly from the ${\boldsymbol{q}}=0$ limit of the above expressions:
@@
$C=\chi_{1}(0,0)=\frac{n_{0}}{\bar{m}}=\frac{1}{V+J},\quad\kappa=\frac{n_{0}}{4\bar{m}^{3}}=\frac{1}{4(V+J)\bar{m}^{2}},\quad\gamma=\frac{n_{0}}{2\bar{m}^{2}}=\frac{1}{2(V+J)\bar{m}}=2\kappa\bar{m}.$ (52)
Above we used the saddle-point condition from Eq. (15) to simplify $C$, $\kappa$ and $\gamma$, leading to the expressions quoted in the main text.
@@ C.1 Magnon spectrum: order of magnitude estimates and numerics
In this Section we estimate the various scales involved in the magnon spectrum obtained from the poles of the magnon propagators $\mathcal{G}_{b\eta}(q)$ in Eq. (27) of the main text. For simplicity we first consider the critical point $\theta_{0}=0$, where the four magnon modes coalesce into only two solutions due to the restored U(1) spin rotation symmetry: a gapped ($+$) and a gapless, linearly dispersing (--) branch:
@@
For rhombohedral multilayers (in particular following spin-orbit proximitized trilayer experiments [44]), we can estimate $J\sim 200$ meV nm2 and polarization densities (per valley) $n_{0}\sim 2\times 10^{11}$ cm-2 relevant for the region where superconductivity is observed. The strength of intra-valley interactions $V$ (which are momentum-independent within our microscopic treatment, Eq. (5)) can be roughly estimated from the value of the gate-screened Coulomb potential
@@
at the Fermi momentum in the relevant low-density regime, $|{\boldsymbol{q}}|\sim k_{F}\sim\left(0.05-0.1\right)a_{0}^{-1}$. Here $q_{e}$ is the electron charge, $\epsilon_{0}$ the vaccum permittivity and $d$ denotes the gate distance, which is of order of a few tens of nanometers in typical experiments, such that $k_{F}d\gg 1$. The estimate for $V$ also depends on the amount of electronic screening in the system, which is difficult to evaluate accurately. In self-consistent Hartree-Fock treatments (e.g., [73, 52, 53, 54]) where screening is accounted for phenomenologically by treating the relative permittivity $\epsilon_{r}$ as a free parameter, a best fit to the experimentally observed phase boundaries [3, 5, 40, 44] is obtained in the range $\epsilon_{r}\sim 20-30$. With these various uncertainties in mind, we estimate $V(|{\boldsymbol{q}}|\sim k_{F})\sim 1-3$ eV nm2. The corresponding ratio of intra-valley to inter-valley interactions in our theory should then be $V/J\sim 5-15$.
@@
To reflect large uncertainties in these estimates, we plot the magnon spectrum in Fig. 3 for $D=1$ and $10$ eV-1, and two values of $V/J=5,10$, both in the ranges justified above. The separation of energy scales between the high-energy and the soft magnon modes is quite large for momentum transfers of the order $k_{0}a_{0}\sim 0.01-0.05$, which are expected to be most important for interband pairing that scatter electrons between majority and minority Fermi surfaces. (Note that $k_{0}$ denotes the momentum transfer between the majority and minority Fermi pockets, which can be much smaller than $k_{F}$ especially in the presence of trigonal warping.)
Using these numbers, we now estimate the scales $\Omega_{0}$ and $g_{\rm eff}$ appearing in our rough sketch of the critical temperature $T_{c}$ in the main text, Eq. (46). For that we focus on the optimal canting angle $\theta_{0}=0$. From Fig. 3(a) and (b), we identify the scale $\Omega_{0}$ with the maximal energy of magnons within the typical momentum transfer scale shown by the shaded grey area. Taking into account the uncertainty in various parameters we take $\Omega_{0}\sim 0.5-5$ meV. The dimensionless coupling constant
$g_{\rm eff}\sqrt{\nu_{+}\nu_{-}}=\frac{2\sqrt{\nu_{+}\nu_{-}}}{Dk_{0}^{2}}$ (69)
is more difficult to estimate as it depends on non-universal band structure quantities. Indeed, both the joint density of states of the majority and minority bands, $\sqrt{\nu_{+}\nu_{-}}$, and the quantity $Dk_{0}^{2}$ which derives from the transverse spin susceptibility, depend crucially on band structure details including trigonal warping and renormalization by interactions. In principle however, this quantity can reach the strong-coupling limit, $g_{\rm eff}\sim O(1)$, when the finite-momentum contribution to the susceptibility $Dk_{0}^{2}$ is comparable to (or smaller than) the density of states.
@@ Appendix D Derivation of magnon-mediated interaction
$\mathcal{M}_{\xi_{1}\xi_{2}}=\frac{1}{2}\frac{1}{Q^{\theta}_{s}Q^{\phi}_{s}+\gamma^{2}\nu^{2}}\left(\begin{matrix}Q^{\phi}_{s}&-\gamma\nu\\ +\gamma\nu&Q^{\theta}_{s}\end{matrix}\right)+\frac{1}{2}\tau_{\xi_{1}}\tau_{\xi_{2}}\frac{1}{Q^{\theta}_{a}Q^{\phi}_{a}+\gamma^{2}\nu^{2}}\left(\begin{matrix}Q^{\phi}_{a}&-\gamma\nu\\ +\gamma\nu&Q^{\theta}_{a}\end{matrix}\right)$ (74)
@@
$g^{+-,+-}_{\xi_{1},\xi_{2}}(q)$ $=-\frac{1}{2}\frac{Q^{\phi}_{s}+(-i)(-\gamma\nu)+(-i)(\gamma\nu)+(-i)^{2}Q^{\theta}_{s}}{Q^{\theta}_{s}Q^{\phi}_{s}+\gamma^{2}\nu^{2}}-\tau_{\xi_{1}}\tau_{\xi_{2}}\frac{1}{2}\frac{Q^{\phi}_{a}+(-i)(-\gamma\nu)+(-i)(\gamma\nu)+(-i)^{2}Q^{\theta}_{a}}{Q^{\theta}_{a}Q^{\phi}_{a}+\gamma^{2}\nu^{2}},$
$=-\frac{1}{2}\frac{Q^{\phi}_{s}-Q^{\theta}_{s}}{Q^{\theta}_{s}Q^{\phi}_{s}+\gamma^{2}\nu^{2}}-\tau_{\xi_{1}}\tau_{\xi_{2}}\frac{1}{2}\frac{Q^{\phi}_{a}-Q^{\theta}_{a}}{Q^{\theta}_{a}Q^{\phi}_{a}+\gamma^{2}\nu^{2}},$ (75)
$g^{-+,+-}_{\xi_{1},\xi_{2}}(q)$ $=-\frac{1}{2}\frac{Q^{\phi}_{s}+(-i)(-\gamma\nu)+i(\gamma\nu)+i(-i)Q^{\theta}_{s}}{Q^{\theta}_{s}Q^{\phi}_{s}+\gamma^{2}\nu^{2}}-\tau_{\xi_{1}}\tau_{\xi_{2}}\frac{1}{2}\frac{Q^{\phi}_{a}+(-i)(-\gamma\nu)+i(\gamma\nu)+i(-i)Q^{\theta}_{a}}{Q^{\theta}_{a}Q^{\phi}_{a}+\gamma^{2}\nu^{2}},$
$=-\frac{1}{2}\frac{Q^{\phi}_{s}+Q^{\theta}_{s}+2i\gamma\nu}{Q^{\theta}_{s}Q^{\phi}_{s}+\gamma^{2}\nu^{2}}-\tau_{\xi_{1}}\tau_{\xi_{2}}\frac{1}{2}\frac{Q^{\phi}_{a}+Q^{\theta}_{a}+2i\gamma\nu}{Q^{\theta}_{a}Q^{\phi}_{a}+\gamma^{2}\nu^{2}}$ (76)
@@
$Q^{\theta}_{s}=2\tilde{J}\cos^{2}\theta_{0}+\kappa\nu^{2}+D{{\boldsymbol{q}}}^{2},\quad Q^{\phi}_{s}=\kappa\nu^{2}+D{{\boldsymbol{q}}}^{2},$ (77)
$Q^{\theta}_{a}=2\tilde{J}\sin^{2}\theta_{0}+\kappa\nu^{2}+D{{\boldsymbol{q}}}^{2},\quad Q^{\phi}_{a}=2\tilde{J}+\kappa\nu^{2}+D{{\boldsymbol{q}}}^{2}.$ (78)
@@
$g^{+-,+-}_{\xi_{1},\xi_{2}}(q)\sim$ $\frac{\tilde{J}\cos^{2}\theta_{0}}{\left[2\tilde{J}\kappa\cos^{2}\theta_{0}+\gamma^{2}\right]\nu^{2}+2\tilde{J}\kappa\cos^{2}\theta_{0}D{{\boldsymbol{q}}}^{2}}$
$-\tau_{\xi_{1}}\tau_{\xi_{2}}\frac{\tilde{J}\cos^{2}\theta_{0}}{4\tilde{J}^{2}\sin^{2}\theta_{0}+\left[2\tilde{J}\kappa\left(1+\sin^{2}\theta_{0}\right)+\gamma^{2}\right]\nu^{2}+2\tilde{J}\kappa(1+\sin^{2}\theta_{0})D{{\boldsymbol{q}}}^{2}},$ (79)
$g^{-+,+-}_{\xi_{1},\xi_{2}}(q)\sim$ $-\frac{\tilde{J}\cos^{2}\theta_{0}+\kappa\nu^{2}+D{{\boldsymbol{q}}}^{2}+i\gamma\nu}{\left[2\tilde{J}\kappa\cos^{2}\theta_{0}+\gamma^{2}\right]\nu^{2}+2\tilde{J}\kappa\cos^{2}\theta_{0}D{{\boldsymbol{q}}}^{2}}$
$-\tau_{\xi_{1}}\tau_{\xi_{2}}\frac{\tilde{J}(1+\sin^{2}\theta_{0})+\kappa\nu^{2}+D{{\boldsymbol{q}}}^{2}+i\gamma\nu}{4\tilde{J}^{2}\sin^{2}\theta_{0}+\left[2\tilde{J}\kappa\left(1+\sin^{2}\theta_{0}\right)+\gamma^{2}\right]\nu^{2}+2\tilde{J}\kappa(1+\sin^{2}\theta_{0})D{{\boldsymbol{q}}}^{2}},$ (80)